Ali spends 2/3 of his money in a shop. a) what fraction of his money remains? b) then he spends 1/4 of the remainder in another shop. What fraction of the original sum will he have left?
1
Expert's answer
2012-12-19T11:15:46-0500
Let& the sum of Ali’s money will be X. So, when he spent 2/3 of X, the rest of money was (X-(2/3)*X) = (1/3)*X.
Then he spent ¼ of (1/3)X, this is (1/12)X
So, the rest of original sum will be x/3-x/12=x/4
Answer: ¼ of original sum